OpenStudy (anonymous):

Given the equation 2x + 4 = 4x - 2, select the reasoning that correctly solves for x. Add 2, subtract 2x, then divide by 2. Add 2, subtract 4x, then divide by -2. Subtract 4, subtract 2x, then divide by -2. Subtract 4, subtract 4x, then divide by 2.

OpenStudy (beccaboo333):

Not fully sure, @ganeshie8 might know better :) he's amazing with math O_O

ganeshie8 (ganeshie8):

to solve x, you need to isolate x

ganeshie8 (ganeshie8):

look at the first statement

OpenStudy (anonymous):

@ganeshie8 are you saying its a?

ganeshie8 (ganeshie8):

2x + 4 = 4x - 2 Add 2, subtract 2x, then divide by 2

ganeshie8 (ganeshie8):

if you add 2 both sides, you wud get : 2x + 6 = 4x

ganeshie8 (ganeshie8):

next, subtract 2x both sides, wat do u get ?

OpenStudy (anonymous):

3

ganeshie8 (ganeshie8):

uhmm no

ganeshie8 (ganeshie8):

try again

ganeshie8 (ganeshie8):

2x + 6 = 4x -2x -2x 6 = 2x

ganeshie8 (ganeshie8):

divide 2 both sides 3 = x

OpenStudy (anonymous):

i meant 2x and then the final answer is 3

ganeshie8 (ganeshie8):

oh you mean final answer is 3

ganeshie8 (ganeshie8):

got you :) you're right then !

OpenStudy (anonymous):

yes @ganeshie8 and thanks you were a big help :)

ganeshie8 (ganeshie8):

So A is your correct option, u can stop here. but if u want u can test the other options same way, and u wil see that they dont make sense :)

OpenStudy (beccaboo333):

Told you he's a math miracle

ganeshie8 (ganeshie8):

very kind of you becca xD wlcme @aleisha96 :)

OpenStudy (anonymous):

thanks and thanks for the medal @ganeshie8